an object of mass 2 kg is moved from infinity to point p .initially that object was at rest but on reaching p its its speed is 2m/s.the work done in moving that object is -4J.Then potential at p is

Grade:12th pass

1 Answers

Rishabh Ranjan
6 years ago
Change in Potential energy = – Work done
U = – W
U = – ( – 4) J = 4 J
Potential = Potential energy / Mass = U/m = 4 / 2 J/kg = 2 J/kg
Therefore answer to this question is 2 J/kg